Plains Wind Energy manufactures wind turbines used in the production of electric power.

The firm is experiencing financial difficulty due to the recent development of competing

sources of energy in the form of extensive shale oil production capacity that is profitable at a

$30 per barrel price for West Texas Intermediate crude. The associated supply-related

decline in the price of energy has reduced demand for Plains’ wind turbines, causing the

value of Plains’ outstanding issue of zero coupon bonds that mature in exactly one year to

drop to $873.17. In the event of a default, bondholders can expect to recover 25 percent of

the bond’s par value of $1000. The annual risk-free rate of interest is 2.5 percent. Assuming

that all interest rates are compounded annually, and that investors value risky bonds by

discounting the expected payoff at the risk-free interest rate:

a. determine the promised yield to maturity for Plains’ outstanding zero-coupon bonds

having one year to maturity, (7 points)

b. estimate the probability that Plains will default on the firm’s outstanding issue of zero

coupon bonds maturing in one year. (8 points)
Answer #1

1.
=1000/873.17-1=14.53%

2.
Let be the probability of default
873.17=(p*25%*1000+(1-p)*1000)/1.025
=>p=0.14